OpenStudy (anonymous):

If f(x) = sqrt x+3 , what is the equation for f^-1(x)

OpenStudy (anonymous):

I only see a file name.

OpenStudy (anonymous):

fixed it

OpenStudy (solomonzelman):

\(\large\color{black}{ y=\sqrt{x+3} }\) right?

OpenStudy (anonymous):

y = sqrtx+3 x = sqrty + 3 y = x^2 - 3 f^1 = x^2 - 3

OpenStudy (anonymous):

yes thats it thank you

OpenStudy (anonymous):

Step 1 : Reorganize your equation to make it y = -------.

OpenStudy (anonymous):

Step 2 : Swap the values of x and y of your equation.

OpenStudy (anonymous):

Isolate for the new value of y.

OpenStudy (anonymous):

Step 4: Replace y by f^-1 =